QUANTUM DYNAMICS STUDIES OF THE VIBRATIONAL STATES OF HO_{3}(X^{2}A′′)

\underline{HUA-GEN YU}, Department of Chemistry, Brookhaven National Laboratory, Upton, NY 11973-5000; BASTIAAN J. BRAAMS, Department of Mathematics and Computer Science, Emory University, Atlanta, Georgia 30322.

We have explored the ground-state potential energy surface for HO_{3} using the HCTH/aug-cc-pVTZ DFT method. Results show that there are two stable isomers, trans-HO_{3} and cis-HO_{3}, with respect to the OH + O_{2} dissociation limit. The binding energies are predicted to be 6.68 kcal/mole (D_{0}) for trans-HO_{3} and 5.62 kcal/mole for cis-HO_{3}. The isomerisation barrier height is only 2.02 kcal/mole from the trans- isomer to the cis- one. The central O-O bond length is obtained as 1.6299 Å in trans-HO_{3} or 1.5958 Å in cis-HO_{3}. A global potential energy surface has been constructed from thousands of ab initio energy points. Based on this surface, we have carried out a full dimensional quantum dynamics study. Results are used to compare with the FTIR experimental observations.^{a}

aThe work at Brookhaven National Laboratory was performed under Contract No. DE-AC02-98CH10886 with the U.S. Department of Energy and supported by its Division of Chemical Sciences, Office of Basic Energy Sciences.